Please help me solve this equation: -1/6(x-18)+1/3(x+3)=x-26

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

Please help me solve this equation: -1/6(x-18)+1/3(x+3)=x-26
---------
Multiply by 6 to eliminate the fractions.

Answer by rothauserc(4718)

-1/6(x-18)+1/3(x+3)=x-26
multiply both sides of = by 6
-(x-18) +2*(x+3) = 6*(x-26)
-x+18 +2x +6 = 6x -156
x +24 = 6x -156
5x = 180
x = 36
check the answer
-1/6*18 +1/3(39) = 10
-3 +13 = 10
10 = 10
answer checks